VARIATIONAL QUANTUM MONTE CARLO CALCULATIONS INCLUDING SPIN-ORBIT COUPLING

\underline{Heinz-Jürgen Flad}, Michael Dolg and Alok Shukla, Max-Planck-Institut für Physik komplexer Systeme (Aussenstelle Stuttgart), D-70569 Stuttgart, Germany.

Electron correlation and relativistic effects play an important role in the chemistry and physics of heavy elements. Relativistic pseudopotentials do not only offer a possibility to reduce the computational effort by means of elimination of the core electrons, but also represent an efficient and reliable way to incorporate scalar relativistic as well as spin-orbit dependent terms in the Hamiltonian. Using this type of valence-only model Hamiltonian we performed variational Quantum Monte-Carlo calculations for atoms and small molecules of heavy elements. Electron correlation was accounted for by Jastrow factors which depend explicitly on the inter-electronic distances. The influence of dynamical correlation on the fine-structure of heavy atoms was investigated. In order to check the accuracy of our valence-only approach, we compare the results to those of fully relativistic all-electron configuration interaction calculations.